OpenStudy (anonymous):

If the legs of a right triangle are 8 and 10, find the length of the hypotenuse.

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Use the Pythagorean theorem: A^2 + B^2 = C^2

OpenStudy (anonymous):

A = 8, B = 10

OpenStudy (anonymous):

how can i put it into a squar root answer?

OpenStudy (anonymous):

So you do this: \[C = \sqrt{a^2 + b^2}\]

OpenStudy (anonymous):

yeah it has too look like this tho

OpenStudy (anonymous):

\[\sqrt{__}\]

OpenStudy (anonymous):